OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). A. y = 7x - 10 B. y = -1/7 x + 2 C. y = 7x - 3 D. y = -7x + 10

OpenStudy (anonymous):

Parallel means the slope is the same, so it will be 7. y+3 = 7(x-1) y+3 = 7x-7 y = 7x-10 so it was A

OpenStudy (anonymous):

@ZhangYan Stop giving answers